The Sun has a radius of 6.9599× 10 8 m. Use the mass of the Sun from your data

booklet to calculate the gravitational field strength at the surface of the Sun.

(274 N/kg)


1
Expert's answer
2022-05-02T13:42:13-0400
"g=G\frac{M}{R^2}\\
=6.67*10^{-11}*\frac{2*10^{30}}{(6.9599*10^8)^2}=275\;\rm m/s^2"